OpenStudy (anonymous):

C= .55t^2 + 550. C= $1430 and t is the number of months. How do I solve this?

OpenStudy (anonymous):

Before you guys comment, I want to thank you for taking the time to help me out. It means a ton.

OpenStudy (anonymous):

It seems like you need to solve for t? \[1430=.55t^2+550\]so move everything to the left side.\[\frac{ 1430-550 }{ .55 }=t^2\]We what we have to get: \[1600=t^2\]Now we take the square root of both sides. \[\sqrt{1600}=\sqrt{t^2}\rightarrow t=40\]

OpenStudy (wolf1728):

I'm guessing this is some kind of interest problem. 1,430 = .55t² + 550 .55t² = 1,430 - 550 .55t² = 880 t² = 1,600 t = 40

OpenStudy (wolf1728):

looks as if we have a consensus on this

OpenStudy (anonymous):

Thanks guys! I wish I can give two medals!

OpenStudy (wolf1728):

thanks grayp